### v5

@@ 1 Introduction
Let $G=(V,E)$ be a simple connected graph on $n\geq 3$ vertices and with $m$ edges. A cut $(S,S^{\prime})$, for any $\emptyset\neq S\subset V$ and $S^{\prime}=V\setminus S$, in $G$ is a partition of its vertices, and the cut-set $\partial S$ is the edges between $S$ and $S^{\prime}$. The (unweighted) edge expansion, also called the Cheeger constant or isoperimetric number or sparsest cut, of $G$ is a ratio that measures the relative number of edges across any vertex partition. It is defined as
$h(G)$ $=\min_{S}\left\{\frac{|\partial S|}{\min\{|S|,|S^{\prime}|\}}:\emptyset\neq S\subset V\right\}$
[…]
where $\partial S=\{(i,j)\in E:i\in S,\,j\in S^{\prime}\}$ is the cut-set associated with any vertex subset $S\subset V$, and $S^{\prime}=V\setminus S$. This constant is positive if and only if the graph is connected, and the exact value tells us that the number of edges across any cut in $G$ is at least $h(G)$ times the number of vertices in the smaller partition. A weighted definition of edge expansion called the conductance of a graph, is
[…]
@@
$\phi(G)$ $=\min_{S}\left\{\frac{|\partial S|}{|S||S^{\prime}|}:\emptyset\neq S\subset V\right\}$
[…]
@@
[…]
@@ 2 Formulations and SDP relaxations
=$\mathcal{F}=\left\{x\in\{0,1\}^{n}:1\leq e^{\top}x\leq\frac{n}{2}\right\}.$
@@
$\begin{split}h(G)&=\min_{x}\;\Big{\{}\dfrac{x^{\top}Lx}{e^{\top}x}:x\in\mathcal{F}\Big{\}}\vskip 6.0pt plus 2.0pt minus 2.0pt\\ &=\min_{x,y}\;\Big{\{}y:\dfrac{x^{\top}Lx}{e^{\top}x}\leq y,\,x\in\mathcal{F}\Big{\}}\vskip 6.0pt plus 2.0pt minus 2.0pt\\ &=\min_{x,y}\;\big{\{}y:x^{\top}Lx-y\,e^{\top}x\leq 0,\,x\in\mathcal{F}\big{\}}.\end{split}$ (1)
@@
$\mathcal{F}_{\operatorname{vol}}=\left\{x\in\{0,1\}^{n}:1\leq d^{\top}x\leq m\right\},$
@@
$h_{\operatorname{vol}}(G)=\min_{x,y}\;\big{\{}y:x^{\top}Lx-y\,d^{\top}x\leq 0,\,x\in\mathcal{F}_{\operatorname{vol}}\big{\}}.$
@@ 2.1 Semidefinite relaxations
where $\tilde{X}$ models $xx^{\top}$ and $k=e^{\top}x$. We add the redundant constraint $\langle J,\widetilde{X}\rangle=k^{2}$ and relax $x\in\{0,1\}^{n}$, $\widetilde{X}=xx^{\top}$ to $\widetilde{X}\succeq 0$ to obtain the above SDP relaxation. To eliminate the variable $k$ in the second, equivalent SDP formulation, we scale $X=\frac{1}{k}\widetilde{X}$.
@@
First observe that $\frac{1}{n}I$ is a strictly feasible point of the primal SDP and the optimum value is finite, hence strong duality holds. The dual of the second SDP in (2) is
$\max_{v}\;\Big{\{}v_{1}-\frac{n}{2}v_{2}+v_{3}:W=L-v_{1}I+(v_{2}-v_{3})J\succeq 0,\,v_{2},v_{3}\geq 0\Big{\}}.$
The eigenvalue of $W$ with respect to the eigenvector $e$ is $-v_{1}+n(v_{2}-v_{3})$. The other eigenvalues of $W$ are then $\lambda_{i}(L)-v_{1}$ for $2\leq i\leq n$. Therefore, we can write the dual as
$\max_{v}\;\Big{\{}v_{1}-\frac{n}{2}v_{2}+v_{3}:n(v_{2}-v_{3})\geq v_{1},\ \lambda_{2}(L)\geq v_{1},\ v_{2},v_{3}\geq 0\Big{\}},$
@@ 2.2 Illustrative examples for motivation
The numbers in the table show that some of these bounds are very weak, in particular, if the number of vertices increases. Interestingly, if we divide the edge expansion problem into $\lfloor\frac{n}{2}\rfloor$ many subproblems with fixed denominator (as we did to obtain the numbers in column 6) the lower bound we obtain by taking the minimum over all SDP relaxations for the subproblems seems to be stronger than the other lower bounds presented in Table 1. We will, therefore, take this direction of computing the edge expansion, namely, we will compute upper and lower bounds on the problem with fixed $k$. Using these bounds will allow to exclude a (hopefully) large number of potential sizes $k$ of the smaller partition. This will leave us with computing the maximum cut of a graph with fixed sizes of the partition $k$ and $n-k$ for a few values of $k$ only.
@@ 3 Fixing the size $k$: Bisection problem
If the size $k$ of the smaller set of the partition of an optimum cut is known, the edge expansion problem would result in a scaled bisection problem. That is, we ask for a partition of the vertices into two parts, one of size $k$ and one of size $n-k$, such that the number of edges joining these two sets is minimized. This problem is NP-hard [11] and has the following formulations for any $k\in\{1,2,\dots,\lfloor\frac{n}{2}\rfloor\}$,
@@ 3.3 Transformation to a max-cut problem
$c_{uw}=\begin{cases}4\mu_{k}(n-2k)&\text{ if }u\in V(G)\text{ and }w=v_{0}\\ 4\mu_{k}-1&\text{ if }uw\in E(G)\\ 4\mu_{k}&\text{ if }uw\not\in E(G)\end{cases}$
@@
The formulation of the max-cut problem in ±$\pm 1$ variables additionally requires $x_{v}=1$ to hold. Because of the symmetry of the cut, we can omit this constraint. Due to our choice of the penalty parameter, it holds that on one side of the maximum cut, there are exactly $k+1$ vertices, including vertex $v$. These $k$ vertices on the same side as $v$ are the vertices in the set of size $k$ in the optimum of the bisection problem.
@@ 4.1 Pre-elimination
1 for $k\in\{1,\dots,\lfloor\frac{n}{2}\rfloor\}$ do
@@ 8       $\mathcal{I}\coloneqq\big{\{}k\in\{1,\dots,\lfloor\frac{n}{2}\rfloor\}\colon\ell_{k}<u^{*}\big{\}}$;
We can further reduce the number of candidates for $k$ by computing a tighter lower bound ~$\tilde{\ell}_{k}$ by solving the DNN relaxation 8 with additional cutting planes. In our implementation we do not compute ~$\tilde{\ell}_{k}$ as part of the pre-elimination but use the lower bound obtained from the solution in the root node of the max-cut solver.
@@
Figures 1 and 2 display the bounds associated with four different graphs. For the graph of the grevlex polytope in dimension 7, considering the bounds $u_{k}$ and ~$\tilde{\ell}_{k}$ the only candidates for $k$ where the optimal solution can be found are 12 and 14. For the grevlex polytope in dimension 8, the sizes 17 and 18 remain as the only candidates. Also for a graph associated to a randomly generated 0/1-polytope and to a network graph, about 2/3 of the potential values of $k$ can be excluded already by considering the cheap lower bound $\ell_{k}$.
@@ 5            $u^{*}\leftarrow u_{k}$;
10       initialize the lower bound for max-cut as $\mathtt{offset}-\lceil u^{*}k\rceil$;
11       compute $h_{k}$ using the max-cut solver;
@@ 5 Parametric optimization
$f(x)=\frac{x^{\top}Lx}{e^{\top}x},\quad\text{and}\quad\mathcal{F}=\Big{\{}x\in\{0,1\}^{n}:1\leq e^{\top}x\leq\frac{n}{2}\Big{\}}.$ (10)
@@
$P(\gamma)=\min_{x}\;\{g_{\gamma}(x):x\in\mathcal{F}\},\qquad\gamma\geq 0.$
@@
$P(0)$ $=\min_{x}\big{\{}x^{\top}Lx:x\in\mathcal{F}\big{\}}$
$=\min_{S}\big{\{}\lvert{\partial S}\rvert:\emptyset\neq S\subset V,|S|\leq\nicefrac{{n}}{{2}}\big{\}}$
$=\min_{S}\big{\{}\lvert{\partial S}\rvert:\emptyset\neq S\subset V\big{\}}$
@@
[…]
@@ 5       $i=i+1$;
Let $\gamma^{\prime}$, $\gamma^{\prime\prime}\in$ and $x^{\prime}$, $x^{\prime\prime}\in\mathcal{F}$ be the optimal solutions of $P(\gamma^{\prime})$ and $P(\gamma^{\prime\prime})$, then
@@ 5.1 Solving the parametrized optimization problem
$Q(\gamma)=\min_{x}\{\gamma_{d}x^{\top}Lx-\gamma_{n}e^{\top}x:x\in\mathcal{F}\}$
@@
Two performant parallel state-of-the-art solvers for binary quadratic programs with linear equality constraints are BiqBin [16] and BiqCrunch [23]. BiqBin first transforms the problem into a QUBO in a pre-processing phase and then solves the equivalent max-cut problem with its max-cut solver. To solve $Q(\gamma)$, we make use of the problem specific properties and directly transform it into a max-cut problem ourselves. The first step towards achieving this is to obtain an exact formulation as a QUBO using binary encoding of the slack variables and the penalty parameter suggested in [17, Thm. 15]. To aid our derivation, let us denote the integer $n_{s}$ and vector $v^{n_{s}}\in^{n_{s}+1}$ by
@@
Let =$\mathcal{F}=\{x\in\{0,1\}^{n}:1\leq e^{\top}x\leq\frac{n}{2}\}$ and $v^{n_{s}}\in^{n_{s}+1}$ with $v^{n_{s}}_{i}=2^{i-1}$ and $n_{s}=\lceil\log_{2}(\lfloor\frac{n}{2}\rfloor)\rceil-1$. Then
=$\mathcal{F}=\Big{\{}x\in\{0,1\}^{n}:e^{\top}x-\alpha^{\top}v^{n_{s}}=1,\ e^{\top}x+\beta^{\top}v^{n_{s}}=\Big{\lfloor}\frac{n}{2}\Big{\rfloor},\ \alpha,\beta\in\{0,1\}^{n_{s}+1}\Big{\}}.$
@@
For any $x\in\mathcal{F}$ it holds that $e^{\top}x=1+s=\lfloor\frac{n}{2}\rfloor-t$ for some slack variables $s$ and $t$ with $10\leq s,t\leq\lfloor\frac{n}{2}\rfloor-$ 1. In fact, any upper bound on $s$ and $t$ greater or equal than $\lfloor\frac{n}{2}\rfloor-1$ is fine, since from $e^{\top}x=1+s$ and $s\geq 0$ it follows that $e^{\top}x\geq 1$ and from $e^{\top}x=\lfloor\frac{n}{2}\rfloor-t$ and $t\geq 0$ it follows that $e^{\top}x\leq\lfloor\frac{n}{2}\rfloor$. The smallest possible value for $n_{s}$ is $\lceil\log_{2}(\lfloor\frac{n}{2}\rfloor)\rceil-1$, since this gives an upper bound of $12^{n_{s}+1}-$ 1 on $s$ and $t$. ∎
@@
$c_{uw}=\begin{cases}-2^{i}\sigma&\text{ if }u=v_{\alpha_{i}}\text{ and }w\in V(G)\\ 2^{i}\sigma&\text{ if }u=v_{\beta_{i}}\text{ and }w\in V(G)\\ 2^{i+j}\sigma&\text{ if }u=v_{\alpha_{i}}\text{ and }w=v_{\alpha_{j}}\\ 2^{i+j}\sigma&\text{ if }u=v_{\beta_{i}}\text{ and }w=v_{\beta_{j}}\\ \end{cases}$

### v6

@@ 1 Introduction
Let $G=(V,E)$ be a simple connected graph on $n\geq 3$ vertices and with $m$ edges. A cut $(S,S^{\prime})$, for any $\emptyset\neq S\subset V$ and $S^{\prime}=V\setminus S$, in $G$ is a partition of its vertices. The (unweighted) edge expansion, also called the Cheeger constant or isoperimetric number or sparsest cut, of $G$ is a ratio that measures the relative number of edges across any vertex partition. It is defined as
$h(G)$ $=\min_{S}\left\{\frac{|\partial S|}{\min\{|S|,|S^{\prime}|\}}\mathrel{\mathop{\ordinarycolon}}\emptyset\neq S\subset V\right\}$
$=\min_{S}\left\{\frac{|\partial S|}{|S|}\mathrel{\mathop{\ordinarycolon}}S\subset V,\,1\leq|S|\leq\frac{n}{2}\right\},$
where $\partial S=\{(i,j)\in E\mathrel{\mathop{\ordinarycolon}}i\in S,\,j\in S^{\prime}\}$ is the cut-set associated with any vertex subset $S\subset V$, and $S^{\prime}=V\setminus S$. This constant is positive if and only if the graph is connected, and the exact value tells us that the number of edges across any cut in $G$ is at least $h(G)$ times the number of vertices in the smaller partition. Another version of edge expansion that accounts for vertex degrees is called the conductance of a graph. It is defined as
$\begin{split}h_{\operatorname{vol}}(G)&=\min_{S}\left\{\frac{|\partial S|}{\min\{\operatorname{vol}(S),\operatorname{vol}(S^{\prime})\}}\mathrel{\mathop{\ordinarycolon}}\emptyset\neq S\subset V\right\}\\ &=\min_{S}\left\{\frac{|\partial S|}{\operatorname{vol}(S)}\mathrel{\mathop{\ordinarycolon}}S\subset V,\,1\leq\operatorname{vol}(S)\leq m\right\},\end{split}$
@@
$\phi(G)$ $=\min_{S}\left\{\frac{|\partial S|}{|S||S^{\prime}|}\mathrel{\mathop{\ordinarycolon}}\emptyset\neq S\subset V\right\}$
$=\min_{S}\,\left\{\frac{|\partial S|}{|S||S^{\prime}|}\mathrel{\mathop{\ordinarycolon}}S\subset V,\,1\leq|S|\leq\frac{n}{2}\right\}.$
@@
For an $n$-vertex graph $G=(V,E)$, the adjacency matrix is a binary matrix $A\in\mathcal{S}^{n}$ having $A_{ij}=1$ if and only if $(i,j)\in E$, and the degree matrix is a $n\times n$ positive diagonal matrix $D$ having $D_{ii}$ equal to the degree of vertex $i\in V$. The Laplacian matrix is L=DAL=D-A, and thus has its nonzero entries as $L_{ii}=\operatorname{deg}(i)$ and $L_{ij}=-1$ for $(i,j)\in E$. The minimum cut in $G$ is defined as $\zeta_{\min}(G)=\min_{\emptyset\neq S\subset V}\,\mathinner{\!\left\lvert\partial S\right\rvert}$.
@@ 2 Formulations and SDP relaxations
$\mathcal{F}=\left\{x\in\{0,1\}^{n}\mathrel{\mathop{\ordinarycolon}}1\leq e^{\top}x\leq\frac{n}{2}\right\}.$
@@
$\begin{split}h(G)&=\min_{x}\;\Big\{\dfrac{x^{\top}Lx}{e^{\top}x}\mathrel{\mathop{\ordinarycolon}}x\in\mathcal{F}\Big\}\vskip 6.0pt plus 2.0pt minus 2.0pt\\ &=\min_{x,y}\;\Big\{y\mathrel{\mathop{\ordinarycolon}}\dfrac{x^{\top}Lx}{e^{\top}x}\leq y,\,x\in\mathcal{F},y\in\mathbb{R}\Big\}\vskip 6.0pt plus 2.0pt minus 2.0pt\\ &=\min_{x,y}\;\big\{y\mathrel{\mathop{\ordinarycolon}}x^{\top}Lx-y\,e^{\top}x\leq 0,\,x\in\mathcal{F},y\in\mathbb{R}\big\}.\end{split}$ (1)
@@
$\mathcal{F}_{\operatorname{vol}}=\left\{x\in\{0,1\}^{n}\mathrel{\mathop{\ordinarycolon}}1\leq d^{\top}x\leq m\right\},$
@@
$h_{\operatorname{vol}}(G)=\min_{x,y}\;\big\{y\mathrel{\mathop{\ordinarycolon}}x^{\top}Lx-y\,d^{\top}x\leq 0,\,x\in\mathcal{F}_{\operatorname{vol}}\big\}.$
@@ 2.1 Semidefinite relaxations
where $\tilde{X}$ models $xx^{\top}$ and $k=e^{\top}x$. We add the redundant constraint $\langle J,\widetilde{X}\rangle=k^{2}$ and relax $x\in\{0,1\}^{n}$, $\widetilde{X}=xx^{\top}$ to $\widetilde{X}\succeq 0$ to obtain the above SDP relaxation. To eliminate the variable $k$ in the second (equivalent) SDP formulation, we scale $X=\frac{1}{k}\widetilde{X}$.
@@
It is well-known (cf. [7, §5.2.6]) that existence of a weak Slater point (i.e., a feasible point satisfying only the nonlinear convex constraints strictly) is sufficient to achieve strong duality for a general convex optimization problem. For the second SDP in (2), the diagonal matrix $\frac{1}{n}I$ is a weak Slater point because it is positive definite and has its trace equal to 1 which is also equal to the sum of all its entries. Hence, strong duality holds with the dual which is
$\max_{v}\;\Big\{v_{1}-\frac{n}{2}v_{2}+v_{3}\mathrel{\mathop{\ordinarycolon}}L-v_{1}I+(v_{2}-v_{3})J\succeq 0,\,v_{2},v_{3}\geq 0\Big\}.$
The psd constraint on the matrix $W=L-v_{1}I+(v_{2}-v_{3})J$ can be simplified after making two observations. First we have $J=ee^{\top}$. Then we deduce, after using the eigendecomposition of the Laplacian $L$, that because $L$ has $e$ as its eigenvector corresponding to the smallest eigenvalue 0, and $I$ is the identity matrix, it follows that $e$ is also an eigenvector of $W$ with eigenvalue $0-v_{1}+e^{\top}e(v_{2}-v_{3})=-v_{1}+n(v_{2}-v_{3})$. The other eigenvalues of $W$ are then $\lambda_{i}(L)-v_{1}$ for $2\leq i\leq n$. Therefore, we can write the dual as
$\max_{v}\;\Big\{v_{1}-\frac{n}{2}v_{2}+v_{3}\mathrel{\mathop{\ordinarycolon}}n(v_{2}-v_{3})\geq v_{1},\ \lambda_{2}(L)\geq v_{1},\ v_{2},v_{3}\geq 0\Big\},$
@@ 2.2 Illustrative examples for motivation
The numbers in the table show that some of these bounds are very weak, in particular, if the number of vertices increases. Interestingly, if we divide the edge expansion problem into $\lfloor\frac{n}{2}\rfloor$ many subproblems with fixed denominator (as we did to obtain the numbers in column 6) the lower bound we obtain by taking the minimum over all SDP relaxations for the subproblems seems to be stronger than the other lower bounds presented in Table 1. We will, therefore, take this direction of computing the edge expansion, namely, we will compute upper and lower bounds on the problem with fixed $k$. Using these bounds will allow to exclude a (hopefully) large number of potential sizes $k$ of the smaller partition. This will leave us with computing the maximum cut of a graph with fixed sizes of the partition $k$ and nkn-k for a few values of $k$ only.
@@ 3 Fixing the size $k$: Bisection problem
If the size $k$ of the smaller set of the partition of an optimum cut is known, the edge expansion problem would result in a scaled bisection problem. That is, we ask for a partition of the vertices into two parts, one of size $k$ and one of size nkn-k, such that the number of edges joining these two sets is minimized. This problem is NP-hard [12] and has the following formulations for any $k\in\{1,2,\dots,\lfloor\frac{n}{2}\rfloor\}$,
@@ 3.3 Transformation to a max-cut problem
[…]
@@
The formulation of the max-cut problem in $\pm 1$ variables additionally requires $x_{v_{0}}=1$ to hold. Because of the symmetry of the cut, we can omit this constraint. Due to our choice of the penalty parameter, it holds that on one side of the maximum cut, there are exactly $k+1$ vertices, including vertex $v_{0}$. These $k$ vertices on the same side as $v_{0}$ are the vertices in the set of size $k$ in the optimum of the bisection problem.
@@ 4.1 Pre-elimination
1for $k\in\{1,\dots,\lfloor\frac{n}{2}\rfloor\}$ do
@@ 8 $\mathcal{I}\coloneqq\big\{k\in\{1,\dots,\lfloor\frac{n}{2}\rfloor\}\colon\ell_{k}<u^{*}\big\}$;
We can further reduce the number of candidates for $k$ by computing a tighter lower bound $\tilde{\ell}_{k}$ by solving the DNN relaxation 8 with additional cutting planes. In our implementation we do not compute $\tilde{\ell}_{k}$ as part of the pre-elimination but use the lower bound obtained from the solution in the root node of the max-cut solver.
@@
Figures 1 and 2 display the bounds associated with four different graphs. For the graph of the grevlex polytope in dimension 7, considering the bounds $u_{k}$ and $\tilde{\ell}_{k}$ the only candidates for $k$ where the optimal solution can be found are 12 and 14. For the grevlex polytope in dimension 8, the sizes 17 and 18 remain as the only candidates. Also for a graph associated to a randomly generated 0/1-polytope and to a network graph, about 2/3 of the potential values of $k$ can be excluded already by considering the cheap lower bound $\ell_{k}$.
@@ 5    $u^{*}\leftarrow u_{k}$;
10   compute $h_{k}$ using the max-cut solver, initialize the lower bound for max-cut as $\mathtt{offset}-\lceil u^{*}k\rceil$;
@@ 5 Parametric optimization
$f(x)=\frac{x^{\top}Lx}{e^{\top}x},\quad\text{and}\quad\mathcal{F}=\Big\{x\in\{0,1\}^{n}\mathrel{\mathop{\ordinarycolon}}1\leq e^{\top}x\leq\frac{n}{2}\Big\}.$ (10)
@@
$P(\gamma)=\min_{x}\;\{g_{\gamma}(x)\mathrel{\mathop{\ordinarycolon}}x\in\mathcal{F}\},\qquad\gamma\geq 0.$
@@
$P(0)$ $=\min_{x}\big\{x^{\top}Lx\mathrel{\mathop{\ordinarycolon}}x\in\mathcal{F}\big\}$
$=\min_{S}\big\{\lvert{\partial S}\rvert\mathrel{\mathop{\ordinarycolon}}\emptyset\neq S\subset V,|S|\leq\nicefrac{{n}}{{2}}\big\}$
$=\min_{S}\big\{\lvert{\partial S}\rvert\mathrel{\mathop{\ordinarycolon}}\emptyset\neq S\subset V\big\}$
@@
where the penultimate equality is from symmetry of the cut function $\mathinner{\!\left\lvert\partial S\right\rvert}=\mathinner{\!\left\lvert\partial S^{\prime}\right\rvert}$. Finiteness of $\mathcal{F}$ and linearity of $g_{\gamma}$ in $\gamma$ tells us that $P$ is the pointwise minimum of finitely many affine (in $\gamma$) functions, and so $P$ is a concave piecewise linear function. The strictly decreasing property was shown in [10, Lemma 3] for general nonlinear fractional problems. ∎
@@ 5 $i=i+1$;
Let $\gamma^{\prime}$, $\gamma^{\prime\prime}\in$ and $^{\prime}, x′′x^{\prime\prime}\in\mathcal{F}$ be the optimal solutions of $P(\gamma^{\prime})$ and $P(\gamma^{\prime\prime})$, then
@@ 5.1 Solving the parametrized optimization problem
$Q(\gamma)=\min_{x}\{\gamma_{d}x^{\top}Lx-\gamma_{n}e^{\top}x\mathrel{\mathop{\ordinarycolon}}x\in\mathcal{F}\}$
@@
Two performant parallel state-of-the-art solvers for binary quadratic programs with linear equality constraints are BiqBin [17] and BiqCrunch [24]. BiqBin first transforms the problem into a QUBO in a pre-processing phase and then solves the equivalent max-cut problem with its max-cut solver. To solve $Q(\gamma)$, we make use of the problem specific properties and directly transform it into a max-cut problem ourselves. The first step towards achieving this is to obtain an exact formulation as a QUBO using binary encoding of the slack variables and the penalty parameter suggested in [18, Thm. 15]. To aid our derivation, let us denote the integer $n_{s}$ and vector $v^{n_{s}}\in^{n_{s}+1}$ by
@@
Let $\mathcal{F}=\{x\in\{0,1\}^{n}\mathrel{\mathop{\ordinarycolon}}1\leq e^{\top}x\leq\frac{n}{2}\}$ and $v^{n_{s}}\in^{n_{s}+1}$ with $v^{n_{s}}_{i}=2^{i-1}$ and $n_{s}=\lceil\log_{2}(\lfloor\frac{n}{2}\rfloor)\rceil-1$. Then
$\mathcal{F}=\Big\{x\in\{0,1\}^{n}\mathrel{\mathop{\ordinarycolon}}e^{\top}x-\alpha^{\top}v^{n_{s}}=1,\ e^{\top}x+\beta^{\top}v^{n_{s}}=\Big\lfloor\frac{n}{2}\Big\rfloor,\ \alpha,\beta\in\{0,1\}^{n_{s}+1}\Big\}.$
@@
For any $x\in\mathcal{F}$ it holds that $e^{\top}x=1+s=\lfloor\frac{n}{2}\rfloor-t$ for some slack variables $s$ and $t$ with $0\leq s,t\leq\lfloor\frac{n}{2}\rfloor-1$. In fact, any upper bound on $s$ and $t$ greater or equal than $\lfloor\frac{n}{2}\rfloor-1$ is fine, since from $e^{\top}x=1+s$ and $s\geq 0$ it follows that $e^{\top}x\geq 1$ and from $e^{\top}x=\lfloor\frac{n}{2}\rfloor-t$ and $t\geq 0$ it follows that $e^{\top}x\leq\lfloor\frac{n}{2}\rfloor$. The smallest possible value for $n_{s}$ is $\lceil\log_{2}(\lfloor\frac{n}{2}\rfloor)\rceil-1$, since this gives an upper bound of $2^{n_{s}+1}-1$ on $s$ and $t$. ∎
@@
[…]